Use the following contribution margin statement for 20X9:

20X9
Sales volume (#units) 100
Revenue $7,500
  Variable costs $4,500
Contribution margin $3,000
  Fixed costs $1,800
Profit $1,200


Required:
a) How much is the price per unit, unit variable cost and unit contribution margin?
price=    unit VC=    unit CM=

b) Write down the CVP relation: profit as a function of sales volume in units
(fill in the missing numbers in an equation like: Profit = 2 *Volume - 50).
Profit =  *Volume -

c) If sales volume increases by 20% (from 100 to 120), how much is the $ change in profits?


d) What is the sales volume required to achieve target profit of $2,100?


e) How much is the breakeven volume?

Breakeven revenue?


f) How much is the margin of safety percentage (at current sales volume of 100 units)?
(enter percentages as a fraction of 1, i.e., enter 23.47% as 0.2347)

g) Based on the margin of safety computed in (f), will you start making a loss if sales drop by 30%?
(enter 1 for yes, 2 for no)

h) How much is the operating leverage (at current sales volume of 100 units)?
(enter percentages as a fraction of 1, i.e., enter 23.47% as 0.2347)

If fixed costs increase, will it increase or decrease the operating risk?
(enter 1=increase, 2=decrease)
Answer #1
Statement showing Computations
Particulars Amount -100 Unist
Sales              7,500.00
Less Variable Expenses            (4,500.00)
Contribution Margin = Sales - VC              3,000.00
Less Fixed cost            (1,800.00)
Net operating income              1,200.00
a) price per unit = 7500/100                    75.00
unit VC =4500/100                    45.00
unit CM =3000/100                    30.00
b) Profit = 30*Volume - 1800
c)  
Contribution =120*30              3,600.00
Less Fixed cost            (1,800.00)
Net operating income              1,800.00
$ change in profits = 1800 - 1200                  600.00
d)
Target Profit              2,100.00
Add Fixed cost              1,800.00
Target contribution              3,900.00
sales volume required to achieve target profit of $2,100 = 3900/30                  130.00
e) Break even Volume =1800/30                    60.00